OpenStudy (anonymous):

Choose the correct simplification of the expression (5a^7b)^2. 25a^14b^2 25a^9b^2 25a^14b 10a^14b^2

OpenStudy (anonymous):

\[5^2a^{7\times 2}b^2\] is a start

OpenStudy (anonymous):

well i dont know how to finish it. i dont think re writing the main equation is going to help.

OpenStudy (anonymous):

i changed a few things, for example now we have \(5^2\) and \(5^2=25\)

OpenStudy (anonymous):

also we have \(7\times 2\) and \(7\times 2=14\)

OpenStudy (anonymous):

oh okay. i didnt see that. i still dont have a clue how to do exponents.

OpenStudy (anonymous):

so the entire expression becomes \[25a^{14}b^2\]

OpenStudy (anonymous):

oh okay. uhm not entirely sure how u came to that.

OpenStudy (anonymous):

i raised everything inside the parentheses to the power of two

OpenStudy (anonymous):

and also used \((a^7)^2=a^{7\times 2}=a^{14}\)

OpenStudy (anonymous):

oh okay thanks !